OpenStudy (rahulmr):

if p=\[\log_{a}2 \], q=\[\log_{a}3 \], r=\[\log_{a}5 \]. Find \[\log_{a}(\frac{ 5\sqrt{3}}{ 3})\] in terms of p, q, r. Anyone help me please. Don't know where to start. Thanks.

OpenStudy (ilcapitano):

When you add two logarithms with the same base together, you multiply their insides: \[\log_{a}2 + \log_{a}3 = \log_{a}(2*3) = \log_{a}6\] Also when you multiply a number against a logarithm, you put the inside to that power! Behold: \[3*\log_{a}2 = \log_{a}(2^3) = \log_{a}8\] ...Yet roots are fractional exponents, right? so... \[(1/2)*\log_{a}3 = \log_{a}(3^{(1/2)}) = \log_{a}(\sqrt{3})\] So these rules come in handy. Can you put them together to come up with your answer?

OpenStudy (rahulmr):

\[\log_{a}5+\frac{ 1 }{ 2 } \log_{a}3-\log_{a}2\] so \[r+\frac{ 1 }{ 2 }q-p\] Should I write it that?

OpenStudy (triciaal):